OpenStudy (anonymous):

The interest that Cindy earns on her investment is given by the equation l = 42t where t is the time in years that her money is invested. She makes a table to determine how much money she would earn for different values of t. She selects 5 as a value for t. What is the corresponding table entry for l? A. 8.4 B. 425 C. 210 D. 420

OpenStudy (texaschic101):

just sub in 5 for t and what do you get ?

OpenStudy (anonymous):

C.

OpenStudy (texaschic101):

I = 42t I = 42(5) I = 210 you are correct :)

OpenStudy (anonymous):

:D I have a few more questions though

OpenStudy (texaschic101):

lets see them

OpenStudy (anonymous):

brb got to go for a sec

OpenStudy (anonymous):

back lol

OpenStudy (anonymous):

The interest that Paula earns on her investment is given by the equation l = 36t where t is the time in years that her money is invested. She makes a table to determine how much money she would earn for different values of t. She selects 12 as a value for t. What is the corresponding table entry for l? A. 3 B. 3612 C. 432 D. 48

OpenStudy (texaschic101):

that was fast

OpenStudy (anonymous):

I thought I posted the message when I left and when I came back I realized I never sent it lol

OpenStudy (texaschic101):

sub in 12 for t

OpenStudy (anonymous):

thought so.

OpenStudy (anonymous):

gtg brb

OpenStudy (texaschic101):

I = 36t I = 36(12) I = you can do this

OpenStudy (anonymous):

I got the nswer @texaschic101 :D thx I have another queston though

OpenStudy (anonymous):

Which is an equation of the given line in standard form? A. –7x + 6y =–11 B. –6x + 7y = 11 C. –6x + 5y = 13 D. –6x +7y =–11

OpenStudy (anonymous):

u there?

OpenStudy (texaschic101):

(1/2, 2) x1 = 1/2 and y1 = 2 (-3,-1) x2 = -3 and y2 = -1 slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (-1-2) / (-3- 0.5) slope(m) = -3/-3.5 slope(m) = 3/3.5 y = mx + b m = 3/3.5 (-3,-1) x = -3 and y = -1 -1 = 3/3.5(-3) + b -1 = -9/3.5 + b -1 + 9/3.5 = b -3.5/3.5 + 9/3.5 = b 5.5 = b y = 3.5x + 5.5 (multiply by 10) 10y = 35x + 55 -35x + 10y = 55 (reduce) -7x + 2y = 11 crap...I can't figure out how they are getting 6 Let me try something else....hold on

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

U THERE?

OpenStudy (anonymous):

sorry caps

OpenStudy (texaschic101):

-6x + 7y = 11 7y = 6x + 11 y = 6/7x + 11/7 y intercept = 11/7 or 1 4/11 x intercept = -6x + 7(0) = 11 -6x = 11 x = - 11/6 or - 1 5/6 Its gonna be this one...answer is B sorry it took so long...I couldn't seem to get my answer to match your answer choices...I finally figured it out

OpenStudy (anonymous):

ohhhh now I get it thx u r awesome :D, last question

OpenStudy (anonymous):

Which is the equation of the given line? A. y = –2 B. x = –2 C. x = 2 D. y = –2x

OpenStudy (texaschic101):

you got a horizontal line....its gonna be y = -2

OpenStudy (anonymous):

thx!

OpenStudy (texaschic101):

was a little faster on this one...lol

OpenStudy (anonymous):

yea XD

OpenStudy (anonymous):

sorry I got one more it turns out plz! @texaschic101

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html ,check here too